Derive integrated rate law for the first order reactions.
1
Expert's answer
2020-01-20T05:11:23-0500

A→ B

v=K1C=-dC/dt

"\int"K1C=-"\int"dC/dt

lnC=-kt+A

if t=0 then C=C0, A=lnC0

lnC=-kt+lnC0

lnC-lnC0=-kt

lnC0-lnC=kt

ln(C0/C)=kt